13. In the current year, the manager of the Downtowner Restaurant has been following the operating budget reduced here:

For the coming year, the following changes are expected:

Given these anticipated changes, prepare ab operating budget for the Downtowner Restaurant for the coming year.

13. In the current year, the manager of the Downtowner Restaurant has been following the operating budget reproduced here For the coming year, the following changes are expected a. Food sales will increase by 10 percent b. Beverage sales will increase by 6 percent C. Food cost percent and beverage cost percent will remain the same d. Fixed salaries and wages $69,300 for this year-will increase by $8,000. Variable salaries and wages will be 16 percent of expected food sales. e. Employee benefits will remain the same percentage of salaries and wages E Controllable expenses will increase by $12,000.Sales Food Beverage Total sales $630,000 140,000 5770,000 Cost of Sales Food 252,000 35,000 Beverages Total costs 287,000 牙8305 Gross Profit Controllable Expenses Salaries and wages Employee benefits Other controllable $173,250 45,045 82,000 expenses Total Controllable Expenses 300,295 182,705 Income before Occupancy and Income Taxes Income before Interest, Costs, Interest, Depreciation, Occupancy Costs 64,000 Depreciation, and Income Taxes Interest Depreciation Total Restaurant Profit 10,000 28,500 38,500 80,205 g Occupancy costs will increase by $5,000. h. Interest and depreciation will remain the same Given these anticipated changes, prepare an operating budget for the Downtowner Restaurant for the coming year
Answer #1

Step 1 Firstly, we need to calculate the required percentages from the current year operating budget which is shown as follows:-

Percentage of Food cost over its sales = (Food Cost/Food Sales)*100

= ($252,000/$630,000)*100 = 40%

Percentage of Beverages cost over its sales = (Beverage cost/Beverage Sales)*100

= ($35,000/$140,000)*100 = 25%

Employee benefits % over salaries & wages in current year = Employee Benefits/Salaries & Wages

= $45,045/$173,250 = 0.26 or 26%

Step 2 Operating budget for the Downtowner Restaurant for the coming year is shown as follows:-

Downtowner Restaurant

Operating Budget for Coming Year (Amounts in $)

Sales
Food (630,000*1.10) 693,000
Beverages (140,000*1.06) 148,400
Total Sales (A) 841,400
Cost of Sales
Food (693,000*40%) 277,200
Beverages (148,400*25%) 37,100
Total costs (B) (314,300)
Gross Profit (C = A-B) 527,100
Controllable Expenses
Salaries and Wages [69,300+8,000+(693,000*16%)] 188,180
Employee Benefits (188,180*26%) 48,927
Other Controllable expenses (Bal fig) (312,295-188,180-48,927) 75,188
Total Controllable expenses (300,295+12,000) (D) (312,295)
Income before occupancy costs, Interest, Dep. and income taxes (E = C-D) 214,805
Less: Occupancy Costs (64,000+5,000) (69,000)
Income before Interest, Dep. and income taxes 145,805
Interest 10,000
Depreciation 28,500
Total (38,500)
Restaurant Profit 107,305

Working Notes:-

1) Total controllable expenses increase by 12,000 from 300,295 to 312,295, therefore other controllable expenses will be equal to total controllable expenses minus expected salaries and wages and expected employee benefits.

2) Expected variable salaries and wages is 16% of expected food sales (i.e. 693,000*16%).